Someone monitored a reaction by observing the time it took for the reactant concentration halve. It was observed that from the initial concentration of 0.01M to 0.005M, it took 10sec, and then from 0.005M to 0.0025M it took 20sec. What reaction order this reaction likely is? How do you make conclusion, please explain.
Answer #1

Concentration of reactant changes from 0.01 M to 0.005 M, it is first half -life.

First half -life (t1/2)1 = 10 sec

similarly , Second- half life (t1/2)2 = 20 sec

half-life(t1/2) x (1/a^(n-1))

a = initial concentration of reactant

(t1/2)1/(t1/2)2 = (a2/a1)^(n-1)

a1 = 0.01 M, a2 = 0.005 M

(10/20) = (0.005/0.01)^(n-1)

n = order of reaction = 2